OpenStudy (anonymous):

The surface areas of two similar solids are 289 ft^2 and 900 ft^2 . the volume of the larger solid id 1755ft^3.what is the volume of the smaller solid?

OpenStudy (anonymous):

ahh here we go again. take the original ratio \[\frac{289}{900}\] take the square root to get \[\frac{17}{30}\] and then cube it to get your answer

OpenStudy (anonymous):

the ratio of the volumes is \[\frac{4913}{27000}\] so multiply by 1755m by that to get smaller volume

OpenStudy (anonymous):

so it would be 319 right

OpenStudy (anonymous):

i didn't do it but i will

OpenStudy (anonymous):

yeah 319 and a little

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yw hope method is clear